# High-order Fusion Graph Contrastive Learning for Recommendation

### v2

@@ Abstract.
Self-supervised learning (SSL) has recently attracted significant attention in the field of recommender systems. Contrastive learning (CL) stands out as a major SSL paradigm due to its robust ability to generate self-supervised signals. Mainstream graph contrastive learning (GCL)-based methods typically implement CL by creating contrastive views through various data augmentation techniques. Despite these methods are effective, we argue that there still exist several challenges. i) Data augmentation (e.g., discarding edges or adding noise) necessitates additional graph convolution (GCN) or modeling operations, which are highly time-consuming and potentially harm the embedding quality. ii) Existing CL-based methods use traditional CL objectives to capture self-supervised signals. However, few studies have explored obtaining CL objectives from more perspectives and have attempted to fuse the varying signals from these CL objectives to enhance recommendation performance.
To overcome these challenges, we propose a High-order Fusion Graph Contrastive Learning (HFGCL) framework for recommendation. Specifically, instead of facilitating data augmentations, we use high-order information from GCN process to create contrastive views. Additionally, to integrate self-supervised signals from various CL objectives, we propose an advanced CL objective. By ensuring that positive pairs are distanced from negative samples derived from both contrastive views, we effectively fuse self-supervised signals from distinct CL objectives, thereby enhancing the mutual information between positive pairs. Experimental results on three public datasets demonstrate the superior recommendation performance and efficiency of HFGCL compared to the state-of-the-art baselines.
@@ 1. INTRODUCTION
Self-supervised learning (SSL) (Liu et al., 2023c, a) has gained increasing recognition for its effectiveness in addressing data sparsity (Xia et al., 2023) challenges. The capability of SSL to extract self-supervised signals from large volumes of unlabeled data allows the approach to compensate for missing information, leading to widespread adoption in numerous studies (Assran et al., 2023; Chen et al., 2023; Baevski et al., 2023). Among various SSL paradigms, contrastive learning (CL) (Jing et al., 2023) stands out by acquiring self-supervised signals through maximizing mutual information between positive pairs in contrastive views. In recent years, the success of graph contrastive learning (GCL)-based methods (Wu et al., 2021; Zhang et al., 2024a) have gained significant attention in the field of recommender systems (Gao et al., 2022a).
@@
In general, for CL to be effective, GCL-based methods require at least two distinct contrastive views. Inspired by various domains (Bayer et al., 2022; Rebuffi et al., 2021), most existing methods employ data augmentation techniques to generate these views. As illustrated in Fig. 1, data augmentation techniques can be broadly classified into two categories: graph augmentation (Wu et al., 2021) and representation augmentation (Yu et al., 2022b). Graph augmentation involves creating interaction subgraphs by randomly discarding nodes or edges on the user-item interaction graph and then generating different contrastive views through graph convolution operations. Representation augmentation generates different contrastive views by adding noise to the embeddings during the graph convolution process. With the generated contrastive views, these methods effectively mine users’ deep preferences, thereby providing more personalized recommendations.
@@
Despite the necessity of generating contrastive views, data augmentation techniques and multiple CL objectives pose several drawbacks. On the one hand. to create diverse contrastive views, data augmentation generally requires additional graph convolution and modeling operations, which significantly increase training cost per epoch. Furthermore, techniques such as graph and representation augmentation, often implemented by discarding edges or adding noise, would potentially harm the quality of embedding between positive pairs, thus affecting the acquisition of self-supervised signals. On the other hand. existing CL-based methods construct CL objectives (Yu et al., 2022b; Yang et al., 2023b) from traditional contrastive views (e.g., user-based and item-based). While the generated self-supervised signals can improve performance to some extent, the fusion of these signals remains problematic, leading to suboptimal recommendation. However, few studies have explored contrastive views from different perspectives and attempted to fuse diverse signals from more CL objectives to further improve recommendation quality.
Based on above analysis, we present two major challenges:
@@
  • How to efficiently obtain high-quality contrastive views without data augmentation?
@@
Figure 1. GCL-based methods using data augmentation. Top: Graph augmentation; Bottom: Representation augmentation.
To address these challenges, we propose a High-order Fusion Graph Contrastive Learning (HFGCL) framework for recommendation. Instead of relying on data augmentations, we derive user-item contrastive views directly from user-item interactions, and demonstrate that the graph convolution networks (GCN) encoder (He et al., 2020) enhances the similarity between user-item pairs. Through analysis in Section 3.2, we find that low-order information, having undergone minimal or no neighborhood aggregation, results in embeddings saturated with self-information. The self-information suppresses the similarity between positive pairs and complicates the maximization of mutual information. Consequently, we exclude low-order embeddings and propose high-order contrastive views. Since the negative samples provided by high-order contrastive views differ from traditional ones (e.g., user views’ negative samples are other users), we incorporate these unique negative samples into our original CL loss. Moreover, to effectively integrate self-supervised signals from different CL objectives, we present a fusion CL loss. By concurrently ensuring that positive pairs remain distant from more negative samples, we further enhance the mutual information between the positive pairs, thereby enabling the effective fusion of diverse self-supervised signals. Overall, HFGCL is a simple yet efficient recommendation framework, capable of fusing rich self-supervised signals from high-order contrastive views with a fusion CL loss, thereby significantly improving recommendation performance. The main contributions of this paper are as follows:
  • We reveal that most GCL-based methods are both time-consuming and potentially detrimental to embedding quality. In addition, We find that existing CL-based methods struggle to effectively fuse self-supervised signals from diverse CL objectives.
  • We propose High-order Fusion Graph Contrastive Learning (HFGCL) for recommendation. HFGCL directly employs high-order information to generate contrastive views and presents a fusion CL loss to integrate different self-supervised signals.
  • We conduct various experimental studies on three publicly datasets, and the results show that HFGCL has significant advantages in terms of recommendation performance and model training efficiency compared to existing state-of-the-art recommendation methods.
@@ 2.1. Graph Neural Networks for Recommendation
Given $\mathcal{U}(|\mathcal{U}|=M)$ and $\mathcal{I}(|\mathcal{I}|=N)$ denote the set of users and items, respectively. In recommender system (Wu et al., 2023), Graph Convolutional Networks (GCN) (Wang et al., 2019; Wu et al., 2019; He et al., 2020) are increasingly favored, which through graph convolution operations, effectively gather high-order neighborhood information to accurately capture user preference behaviors. LightGCN (He et al., 2020) as currently the most popular GCN encoder, which effectively captures high-order information through neighborhood aggregation. For efficient training, graph Laplace matrix is proposed: $\tilde{A}=D^{-\frac{1}{2}}A{D^{-\frac{1}{2}}}$, where $\tilde{A}\in\mathbb{R}^{(M+N)\times(M+N)}$ denotes graph Laplacian matrix, and $A$ denotes adjacency matrix, and $D$ denotes diagonal matrix of A. To ensure that diverse neighborhood information is integrated, final embeddings are aggregated from all layers: $\mathbf{E}=\frac{1}{L+1}(\mathbf{E}^{(0)}+...+\tilde{A}^{L}\mathbf{E}^{(0)}),$ where $L$ denotes the number of GCN layers. The point-wise Bayesian Personalisation Ranking (BPR) loss (Rendle et al., 2009) is adopted to optimize model parameters:
@@
where $<u,i,j>$ represents a triplet input to the model.
@@ 2.2. Contrastive Learning for Recommendation
Recent studies have demonstrated that contrastive learning (CL), through the generation of self-supervised signals, effectively addresses the challenge of data sparsity in recommender systems (Xia et al., 2023). CL-based methods create contrastive views through data augmentation and seek to optimize the mutual information between positive pairs, thereby obtaining self-supervised signals. Most of the existing methods mainly use InfoNCE (Chen et al., 2020) loss for CL:
(2) $\mathcal{C}(\mathbf{e}_{a},\mathbf{e}_{b})=-\sum_{<a,b>\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathbf{e}^{\top}_{a}\mathbf{e}_{b}/\tau)}{\sum_{c\in\mathcal{B}}\mathrm{exp}(\mathbf{e}^{\top}_{a}\mathbf{e}_{c}/\tau)},$
where $a$ and $b$ denote samples ($e.g.,$ user or item) in same batch $\mathcal{B}$, $\tau$ denotes the temperature coefficient. CL loss aims to maximizes the mutual information between positive pairs $\mathbf{e}_{a}$ and $\mathbf{e}_{b}$, and to increase the distance between $\mathbf{e}_{a}$ and other negative sample $\mathbf{e}_{c}$.
@@ 2.3. Effectiveness of Data Augmentation
Existing researches on CL underscore the critical role of data augmentations (Wu et al., 2021; Yu et al., 2022b; Yang et al., 2023b; Zhang et al., 2024b), which is attributed to CL that necessitates the creation of diverse contrastive views for effective implementation. In recommendation, there are two main types in Fig. 1: (1) graph augmentation (Wu et al., 2021) (2) representation augmentation (Yu et al., 2022b).
@@
While these contrastive views are effective, data augmentations corrupt the embedding’s quality. In Fig. 2, graph augmentation methods discard edges when generating contrastive views, leading to relatively poorer embedding feature distributions and density distributions. Representation augmentation methods by adding noise to the graph convolution. Although the effect of some noise on the distribution is less pronounced than that of graph augmentation, it still results in less smooth density distribution. Additionally, data augmentations are exceptionally time-consuming. Table 1 provides the per-epoch time consumption for the GCN-based method ($e.g.,$ LightGCN) and GCL-based methods with ($e.g.,$ SGL and SimGCL) and without ($e.g.,$ HFGCL) data augmentations. SGL (Wu et al., 2021), which utilizes the graph augmentation technique to generate subgraphs, is extremely time-consuming. Additionally, most GCL-based ($e.g.,$ SimGCL and VGCL) methods require repetitive graph convolution or modeling operations to obtain the contrastive views, further contributing to the time consumption. This raises the question: Are data augmentation techniques necessary for contrastive learning in recommendation? In fact, despite the demonstrated effectiveness of data augmentations, how to efficiently generate high quality contrastive views remains a key challenge in CL.
Table 1. Efficiency comparison of the GCL-based methods per epoch time consuming across three datasets. (s: second)
LightGCN 76.6 (s) 37.7 (s) 65.2 (s)
SGL 188.1 (s) 92.5 (s) 154.5 (s)
SimGCL 194.5 (s) 83.5 (s) 167.9 (s)
VGCL 98.7 (s) 51.4 (s) 81.7 (s)
HFGCL 78.6 (s) 38.9 (s) 66.2 (s)
(a) Amazon-book
                        Feature
         Density
          Angles
                                Angles
                                                      Angles
(b) Yelp2018
                        Feature
         Density
          Angles
                                Angles
                                                      Angles
Figure 2. Distribution of item representations: We follow previous work (Yu et al., 2022b) to plot feature & density distributions. HFGCL has more uniform feature distributions and smoother density distributions due to use of unperturbed embeddings information.

### v4

# Fusion Self-supervised Learning for Recommendations

@@ Abstract.
Recommender systems are widely deployed in various web environments, and self-supervised learning (SSL) has recently attracted significant attention in this field. Contrastive learning (CL) stands out as a major SSL paradigm due to its robust ability to generate self-supervised signals. Mainstream graph contrastive learning (GCL)-based methods typically implement CL by creating contrastive views through various data augmentation techniques. Despite these methods are effective, we argue that there still exist several challenges. i) Data augmentation ($e.g.,$ discarding edges or adding noise) necessitates additional graph convolution (GCN) or modeling operations, which are highly time-consuming and potentially harm the embedding quality. ii) Existing CL-based methods use traditional CL objectives to capture self-supervised signals. However, few studies have explored obtaining CL objectives from more perspectives and have attempted to fuse the varying signals from these CL objectives to enhance recommendation performance.
To overcome these challenges, we propose a Fusion Self-supervised Contrastive Learning framework for recommendation. Specifically, instead of facilitating data augmentations, we use high-order information from GCN process to create contrastive views. Additionally, to integrate self-supervised signals from various CL objectives, we propose an advanced CL objective. By ensuring that positive pairs are distanced from negative samples derived from both contrastive views, we effectively fuse self-supervised signals from distinct CL objectives, thereby enhancing the mutual information between positive pairs. Experimental results on three public datasets demonstrate the superior recommendation performance and efficiency of HFGCL compared to the state-of-the-art baselines.
@@ 1. INTRODUCTION
Recommender systems are essential to a variety of web platforms (Wu et al., 2023b), including e-commerce and streaming services, where they enhance user experiences by delivering personalized content. However, these systems often face the challenge of data sparsity. Self-supervised learning (SSL) (Liu et al., 2023b, a) has gained increasing recognition for its effectiveness in addressing data sparsity (Xia et al., 2023). The capability of SSL to extract self-supervised signals from large volumes of unlabeled data allows the approach to compensate for missing information, leading to widespread adoption in numerous studies (Assran et al., 2023; Chen et al., 2023; Baevski et al., 2023). Among various SSL paradigms, contrastive learning (CL) (Jing et al., 2023) stands out by acquiring self-supervised signals through maximizing mutual information between positive pairs in contrastive views. In recent years, the success of graph contrastive learning (GCL)-based methods (Wu et al., 2021; Zhang et al., 2024a) have gained significant attention in the field of recommender systems (Gao et al., 2022a).
@@
In general, for CL to be effective, GCL-based methods require at least two distinct contrastive views. Inspired by various domains (Bayer et al., 2022; Rebuffi et al., 2021), most existing methods employ data augmentation techniques to generate these views. As illustrated in Fig. 1, data augmentation techniques can be broadly classified into two categories: graph augmentation (Wu et al., 2021) and feature augmentation (Yu et al., 2022). Graph augmentation involves creating interaction subgraphs by randomly discarding nodes or edges on the user-item interaction graph and then generating different contrastive views through graph convolution operations. Feature augmentation generates different contrastive views by adding noise to the embeddings during the graph convolution process. With the generated contrastive views, these methods effectively mine users’ deep preferences, thereby providing more personalized recommendations.
@@
Despite the necessity of generating contrastive views, data augmentation techniques and multiple CL objectives pose several drawbacks. On the one hand. To create diverse contrastive views, in addition to the neighborhood-aggregated embeddings required for the main recommendation task, data augmentation generally requires additional graph convolution and modeling operations based on the same initial embeddings, which significantly increase training cost per epoch. Furthermore, graph augmentation in data augmentation often randomly drop nodes or edges, thereby disrupting the inherent properties of the input graph. In contrast, feature augmentation alleviate this issue by adding noise to each node, but they still overlook the unique characteristics of individual nodes. On the other hand. Existing GCL-based methods construct CL objectives (Yu et al., 2022; Yang et al., 2023b) from traditional contrastive views ($e.g.,$ user-based and item-based). While the generated self-supervised signals can improve performance to some extent, the fusion of these signals remains problematic, leading to suboptimal recommendation (as analyzed in Section 3.2). In addition, few studies have explored contrastive learning paradigm from different perspectives and attempted to fuse diverse signals from more CL objectives to further improve recommendation quality.
Based on the above analysis, we identify two key challenges critical for advancing this field:
@@
  • How to efficiently obtain high-quality contrastive views without data (graph and feature) augmentations?
@@
Figure 1. GCL-based methods using data augmentation. Top: Graph augmentation; Bottom: Feature augmentation.
To address these challenges, we propose a High-order Fusion Graph Contrastive Learning (HFGCL) framework for recommendation. Instead of relying on data augmentations, we uncover the user-item contrastive paradigm directly from users’ interaction behavior, and demonstrate that the graph convolution (GCN) encoder (He et al., 2020) enhances the similarity between user-item pairs. Through analysis in Section 4.2, we find that low-order information, having undergone minimal or no neighborhood aggregation, results in embeddings saturated with self-information. The self-information suppresses the similarity between positive pairs and complicates the maximization of mutual information. Consequently, we exclude low-order information and propose high-order contrastive views. Since the negative pairs provided by high-order contrastive views differ from traditional ones ($e.g.,$ user views’ negative pairs are other users), we integrate the users and items in the batch and introduce them into the original CL loss. Moreover, to effectively integrate self-supervised signals from different CL objectives ($i.e.,$ user-user, item-item, user-item contrastive views), we present a fusion CL loss. By reconstructing the fused embeddings between users and items, we further enhance the mutual information between the positive pairs, thereby enabling the effective fusion of diverse self-supervised signals. Overall, HFGCL is a simple yet efficient recommendation framework, capable of fusing rich self-supervised signals from multi-scale contrastive objectives with a fusion CL loss, thereby significantly improving recommendation performance.
The main contributions of this paper are as follows:
  • We reveal that most GCL-based methods are time-consuming to generate contrastive views and that existing CL methods struggle to fuse self-supervised signals from diverse contrastive objectives.
  • We introduce a novel graph contrastive learning paradigm from the perspective of more effective contrastive views construction, and propose the High-order Fusion Graph Contrastive Learning (HFGCL) framework for recommendation.
  • We leverage stacking multiple GCN layers to generate high-order contrastive views, and design a multi-scale fusion contrastive learning objective to better integrate self-supervised signals.
  • We conduct various experimental studies on three publicly datasets, and the results show that HFGCL has significant advantages in terms of recommendation performance and method training efficiency compared to existing the-state-of-art baselines.
@@ 2.1. Graph Neural Networks for Recommendation
Given $\mathcal{U}(|\mathcal{U}|=M)$ and $\mathcal{I}(|\mathcal{I}|=N)$ denote the set of users and items, respectively. In recommender system (Wu et al., 2023a), Graph Convolutional Networks (GCN) (Wang et al., 2019; Wu et al., 2019; He et al., 2020) are increasingly favored, which through graph convolution operations, effectively gather high-order neighborhood information to accurately capture user preference behaviors. LightGCN (He et al., 2020) as currently the most popular GCN encoder, which effectively captures high-order information through neighborhood aggregation. For efficient training, graph Laplace matrix is proposed: $\tilde{A}=D^{-\frac{1}{2}}A{D^{-\frac{1}{2}}}$, where $\tilde{A}\in\mathbb{R}^{(M+N)\times(M+N)}$ denotes graph Laplacian matrix, and $A$ denotes adjacency matrix, and $D$ denotes diagonal matrix of $A$. To ensure that diverse neighborhood information is integrated, final embeddings are aggregated from all layers: $\mathbf{E}=\frac{1}{L+1}(\mathbf{E}^{(0)}+...+\tilde{A}^{L}\mathbf{E}^{(0)}),$ where $L$ denotes the number of GCN layers. The point-wise Bayesian Personalisation Ranking (BPR) loss (Rendle et al., 2009) is adopted to optimize model parameters:
@@
where $\langle u,i,j\rangle$ represents a triplet input to the model.
@@ 2.2. Contrastive Learning for Recommendation
Recent studies (Wu et al., 2021) have demonstrated that contrastive learning (CL), the primary paradigm in self-supervised learning, effectively addresses the challenge of data sparsity in recommender systems through the generation of self-supervised signals (Xia et al., 2023). CL-based methods create contrastive views through data augmentation and seek to optimize the mutual information between positive pairs, thereby obtaining self-supervised signals. Most of the existing methods mainly use InfoNCE (Chen et al., 2020) loss for CL:
(2) $\mathcal{C}(\mathbf{e}_{a},\mathbf{e}_{b})=-\sum_{\langle a,b\rangle\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathbf{e}^{\top}_{a}\mathbf{e}_{b}/\tau)}{\sum_{c\in\mathcal{B}_{b}}\mathrm{exp}(\mathbf{e}^{\top}_{a}\mathbf{e}_{c}/\tau)},$
where $\langle a,b\rangle$ denote positive pair, assumed here as $\langle a_{2},b_{3}\rangle$ in same batch =$\mathcal{B}=\{\langle a_{1},b_{2}\rangle,\langle a_{2},b_{3}\rangle,...,\langle a_{m},b_{n}\rangle\}$, $c$ denotes all samples of $\mathcal{B}_{b}=\{b_{2},b_{3},...,b_{n}\}$ in batch $\mathcal{B}$, $\tau$ denotes the temperature coefficient. CL loss aims to maximizes the mutual information between positive pairs $\mathbf{e}_{a}$ and $\mathbf{e}_{b}$, and to increase the distance between $\mathbf{e}_{a}$ and other negative sample $\mathbf{e}_{c}$.
3. EXPLORING GRAPH CONTRASTIVE LEARNING FOR RECOMMENDATION
In this section, we explore the key aspects of GCL-based recommendation methods, with a particular focus on data augmentation techniques and contrastive views.
@@ 3.1. Effectiveness of Data Augmentation
Table 1. Comparison of GCN and GCL-based methods (with Data Augmentations) in terms of per-epoch runtime across three datasets (s: seconds, x: times slower).
LightGCN 67.9s 34.3s 54.4s
SGL-ED 188.1s (2.8x) 92.5s (2.7x) 154.5s (2.8x)
SimGCL 194.5s (2.9x) 83.5s (2.4x) 167.9s (3.1x)
NCL 137.1s (2.0x) 60.5s (1.8x) 89.5s (1.7x)
CGCL 140.5s (2.1x) 55.6s (1.6x) 132.2s (2.4x)
VGCL 98.7 (1.5x) 51.4 (1.5x) 81.7 (1.5x)
Existing researches on CL underscore the critical role of data augmentations (Wu et al., 2021; Yu et al., 2022; Yang et al., 2023b; Zhang et al., 2024b), which is attributed to CL that necessitates the creation of diverse contrastive views for effective implementation. In recommendation, there are two main types in Fig. 1: (1) graph augmentation (Wu et al., 2021) (2) feature augmentation (Yu et al., 2022).
Graph augmentation typically involves generating two distinct subgraphs by randomly discarding edges/nodes from original graph:
@@
where $\mathcal{G}$ denotes the original graph, $\mathcal{N}$ and $\mathcal{E}$ denote the nodes and edges, respectively, $\mathcal{G}^{\prime}$ and $\mathcal{G}^{\prime\prime}$ denotes the augmentation subgraphs, $p$ denotes keeping rate. Then, we perform graph convolution operations by $\mathcal{G}^{\prime}$ and $\mathcal{G}^{\prime\prime}$ to obtain different views $\mathbf{e}^{\prime}$ and $\mathbf{e}^{\prime\prime}$ for the same node $\mathbf{e}$. Unlike graph augmentation, feature augmentation has garnered significant attention in numerous studies by introducing noises into graph convolution process:
(4) $\mathbf{e}^{\prime}=\mathbf{e}+\triangle,\mathbf{e}^{\prime\prime}=\mathbf{e}+\triangle,$
where $\triangle$ denotes the added noise ($e.g.$, gaussian noise or uniform noise). After obtaining the contrastive views, contrastive learning is performed using the CL loss in Eq. 2, $\mathcal{L}_{cl}=\mathcal{C}(\mathbf{e}^{\prime},\mathbf{e}^{\prime\prime})$, effectively extracting self-supervised signals from different views to alleviate the data sparsity problem.
@@
Despite these contrastive views are effective in improving performance (as shown in Table 2), data augmentation has a significant drawback: generating contrastive views each epoch adds extra time, significantly increasing training costs. Table 1 provides the per-epoch time consumption for the GCN-based method ($e.g.,$ LightGCN (He et al., 2020)) and GCL-based methods with ($e.g.,$ SGL-ED (Wu et al., 2021), SimGCL (Yu et al., 2022), NCL (Lin et al., 2022), CGCL (He et al., 2023) and VGCL (Yang et al., 2023b)). For fair comparison, all methods are tested under the same experimental setup, with details provided in Section 5.1. SGL-ED, which utilizes the graph augmentation technique to generate subgraphs, is extremely time-consuming. Additionally, most GCL-based ($e.g.,$ SimGCL) methods require repetitive graph convolution or modeling operations to obtain the contrastive views, further contributing to the high time cost. This raises an important question: Is it possible to propose a novel contrastive learning paradigm that does not rely on data augmentation techniques, but still generates high-quality contrastive views and improves training efficiency? While data augmentations have proven effective in generating contrastive views and enhancing recommendation performance, efficiently generating high-quality contrastive views remains a key challenge in contrastive learning.
3.2. Necessity of More Contrastive Views
Table 2. Contrastive View and Performance of GCN and GCL-based Methods. (User and Item denote the number of User-based and Item-based contrastive views, respectively.)
Method CL Views Amazon-Book Yelp2018
User Item R@20 N@20 R@20 N@20
LightGCN - - 0.0411 0.0315 0.0639 0.0525
SGL-ED 2 2 0.0478 0.0379 0.0675 0.0555
SimGCL 2 2 0.0515 0.0414 0.0721 0.0601
NCL 4 4 0.0481 0.0373 0.0685 0.0577
CGCL 6 6 0.0483 0.0380 0.0690 0.0560
VGCL 4 4 0.0515 0.0410 0.0715 0.0587
[…]
Figure 2. Framework of HFGCL. Left: User-item interaction. Center: High-order contrastive views generation process. In the figure, let $h=2$ and $L=3$. After aggregating high-order embeddings, user $u_{2}$ has a representation more similar to the positive item $i_{2}$, with only $u_{4}$ missing. This indicates a close relationship between $u_{2}$ and $i_{2}$ in the high-order embedding space. Right: Dual optimization objectives. During the model training phase, two optimization objectives are included: the primary and the contrastive optimization objective, refer to Fig. 3 for the meaning.
Our experimental results suggest that while generating more contrastive views can provide richer self-supervised signals, the subtle differences among these signals pose significant challenges for effective utilization. Consequently, methods that prioritize generating a larger number of contrastive views often exhibit relatively inferior performance compared to SimGCL, which with high-quality contrastive views. As discussed in the previous section, identifying and generating high-quality contrastive views remains a critical yet often overlooked challenge in contrastive learning. Furthermore, as suggested by previous studies (Yang et al., 2023b), although self-supervised signals from different views demonstrate potential in alleviating data sparsity, the nuanced discrepancies between these signals significantly hinder their seamless fusion, ultimately limiting the effectiveness of these new methods.